Barry bought 25 stamps of 5c and 20c. If the total value of the stamps is $4.40 how many stamps of each kind did he buy.
Answer by macston(5194)

.
F=number of 5c stamps; T=number of 20c stamps
.
F+T=25
T=25-F
.
5F+20T=440
5F+20(25-F)=440
5F+500-20F=440
-15F=-60
F=4
ANSWER: He bought four 5c stamps.
.
T=25-F=25-4=21
ANSWER 2: He bought twenty one 20c stamps.
.
CHECK:
.
5F+20T=440
5(4)+20(21)=440
20+420=440
440=440